OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (anonymous):

\[u\cdot v=\|u\|\|v\|\cos\theta\] Find the dot product and norms, then solve for the angle.

OpenStudy (anonymous):

So -34costheta @SithsAndGiggles

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

\[\langle6,-1\rangle\cdot\langle7,-4\rangle=6\times7+(-1)\times(-4)=42+4=46\] \[\|\langle6,-1\rangle\|=\sqrt{6^2+(-1)^2}=\sqrt{37}\] \[\|\langle7,-4\rangle\|=\sqrt{7^2+(-4)^2}=\sqrt{65}\] \[\cos\theta=\frac{46}{\sqrt{37}\sqrt{65}}~~\implies~~\theta=\text{you can compute}\]

OpenStudy (anonymous):

.917 @si

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@mathstudent55